# Car-Following Models: A Multidisciplinary Review

### v4

@@ Abstract
Car-following (CF) algorithms are crucial components of traffic simulations and have been integrated into many production vehicles equipped with Advanced Driving Assistance Systems (ADAS). Insights from the model of car-following behavior help us understand the causes of various macro phenomena that arise from interactions between pairs of vehicles. Car-following models encompass multiple disciplines, including traffic engineering, physics, dynamic system control, cognitive science, machine learning, and reinforcement learning. This paper presents an extensive survey that highlights the differences, complementarities, and overlaps among microscopic traffic flow and control models based on their underlying principles and design logic. It reviews representative algorithms, ranging from theory-based kinematic models, Psycho-Physical Models, and Adaptive cruise control models to data-driven algorithms like Reinforcement Learning (RL) and Imitation Learning (IL). The manuscript discusses the strengths and limitations of these models and explores their applications in different contexts. This review synthesizes existing researches across different domains to fill knowledge gaps and offer guidance for future research by identifying the latest trends in car following models and their applications.
@@
Car-Following Behavior, Adaptive Cruise Control, Theory-based, Data-driven.
@@ I Introduction
To furnish a complete picture of traffic flow, vehicles on the road are studied as a whole - not merely as isolated particles. Car-following (CF) is the fundamental driving behavior that forcefully constrains the ego-vehicle from adopting a specific pattern in the presence of other vehicles (Figure 1). Since the 1950s or even earlier, many researchers have focused on driver behavior modeling and applied such models to the simulation of Human Driver Vehicles (HDVs) to investigate micro- and macro-traffic phenomena [1, 2]. The expansion of autonomy and connectivity in the transportation system has escalated the hype for integration of car following models with Connected and Automated Vehicle (CAV) technologies, such as collaborative driving, and Vehicle-to-Everything (V2X) communication. Autonomous Vehicles (AVs) with different parameter settings or objective functions become a new source of variety and add more complexities to the existing roadway systems. The heterogeneity of vehicle types and driver behaviors gives rise to various challenges for traffic operation, which must be comprehensively analyzed before upgrading existing infrastructure. Car-following models become a powerful tool to address challenges mentioned above and answer those critical questions in the upcoming era of the connected and automated transportation systems. By analyzing the interaction between vehicles, researchers can identify efficient driving strategies that includes optimizing speeds, minimizing unnecessary acceleration and deceleration events, and maintaining optimal following distances to ensure smooth traffic flow. Modern adaptive cruise control (ACC) systems already use basic car-following principles to maintain safe distances between vehicles. Decades of theory-based driving behavior research are now merging with data-driven opportunities (see Figure 2). Although the data-driven model will undoubtedly become critically important, the traditional driver behavior expertise remains relevant. It is auspicious to integrate theory-based models with the fast-evolving AI models, which can reason, learn, and fill in missing information to achieve better interpretability, generalizability, and reliability.
@@
Figure 2: The Historical Timeline of Reviewed Car Following Models.
Existing survey papers[3, 4, 5, 6, 7, 8, 11, 9, 10, 12, 13, 14, 15] lack of structured framework and often have limitations in terms of scope and perspective (see Table I). For example, a frequently referenced paper [4] mainly surveyed from psychologic perspective and did not consider Intelligent Driver Model (IDM) and Optimal Velocity Model (OVM). Some surveys [10, 8] reviewed multiscale traffic flow models from the engineering and physics perspective, while many control theory models were omitted. Other literature reviews [7, 9] are limited to theory-based models and inadequate to reflect the data-driven algorithms. By taking into account various domains and angles, this review paper aims to provide a more structured categorization of the car following models (see Figure 3), addressing the potential gaps left by previous surveys with a narrower focus. A multidisciplinary review can be particularly beneficial in fields like intelligent vehicle technology, where insights from different disciplines – such as electrical & computer engineering, mechanic engineering, traffic flow theory, physics, and optimal control – are crucial for developing a complete picture of how these agents operate and interact with their environment. This review paper not only integrates diverse knowledge but also promotes the discovery of innovative solutions that might be overlooked when considering a single perspective. This broader approach allows for a more thorough evaluation of the state-of-the-art, challenges, and future directions of car-following models, making a unique contribution to the field.
@@
Summarized car-following models using mathematical model descriptions, stability analysis, calibration from empirical testing and evaluation and finally automatic vehicular control systems.
It only covers stimulus-response models from psychological and behavior perspective.
Car-following: a historical review [4]
1999
@@
Summarized theoretical early-stage car-following models from psychological and behavior perspective, such as safety distance, desired gaps, stimulus-response process.
It doesn’t include some important models such as Intelligent Driver Model (IDM) and Optical Velocity Model (OVM).
Comparison of Car-following models [5]
2004
@@
Compared theoretical car-following models used in well-known traffic micro-simulation software, including AIMSUN, MITSIM, Paramics and VISSIM.
From a software perspective to evaluate how different programs could replicate real world scenarios.
Driving Behavior: models and challenges [6]
2007
@@
Reviewed and expanded review lists in [4] with IDM, OVM and Cellular Automata. Discussed the challenges associated with the myopic driver behavior models.
Only summarized theory-based car-following models from human behavior perspective.
Microscopic car-following model for the traffic flow: the state of the art [7]
2012
@@
Summarized and reviewed theory-based car-following models, also provided the stability analysis from control theory point of view.
Only summarized theory-based car-following models from human behavior perspective.
Traffic flow dynamics[8]
2013
@@
Categorized existing models as elementary models and strategic models by design principles in microscopic traffic flow modeling.
Only covers kinematic models from the physics perspectives.
Incorporating human-factors in car-following models: a review of recent developments and research needs[9]
2014
Introduced various human factors that could potentially impact driving styles.
@@
Only summarized theory-based car-following models from engineering and human behavior perspective.
Genealogy of traffic flow models [10]
2015
@@
Summarized macro-, meso- and micro- traffic flow models, offered an historical model tree by tracing genealogy of the models.
Reviewed and surveyed car following models from the traffic engineering and physical perspectives
A state-of-the-art review of car-following models with particular considerations of heavy vehicles [11]
2015
Categorized CF models into classical models and AI models. Emphasized the impact of heavy trucks in car following behavior.
@@
Under AI category, only fuzzy logic models and Neural Network models were considered. Under the classical model category, it doesn’t include most recent adaptive control models
Simulation Strategies for Mixed Traffic Conditions: A Review of Car-Following Models and Simulation Frameworks [12]
2020
@@
This review paper focus on Heterogeneous vehicle types and mixed traffic flow modeling
Inherit the same categorization from previous review paper [4] and mainly from simulation software perspective
@@
A review of car-following models and modeling tools for human and autonomous-ready driving behaviors in micro-simulation [13]
2021
Emphasized the differences between human-driven vehicles and autonomous vehicles in car-following models.
Inherit the same categorization from previous review paper [4] and mainly from simulation software perspective.
The Car-Following Model and Its Applications in the V2X Environment: A Historical Review [14]
2021
Emphasize applications of car-following models in the V2X environment
@@
Use the similar categorization framework from previous review paper [4, 6] to categorize car-following models
Modeling the Car-Following Behavior with Consideration of Driver, Vehicle, and Environment Factors: A Historical Review [15]
2022
@@
This review paper surveyed car models that consider diverse driver styles, physical and dynamic characteristics of vehicles, as well as environmental factors.
Inherit the similar categorization from previous review paper [4, 6, 14] and is developed from traffic engineering perspective.
@@
Figure 3: Proposed Taxonomy of Car Following Models and Representative Algorithms. Kinematic Models are Divided into First-Order and Second-Order Subcategories.
@@ II-A1 First-Order Kinematic Models
In the formulation, $a_{n}$ is the maximum acceleration; $b_{n}$ is the most severe braking that the driver of vehicle $n$ wishes to undertake; $v_{n}$ is the speed at which the driver of vehicle $n$ wishes to travel; $V_{n}$ is the desired speed of vehicle $n$; $\tau$ is the reaction time; $\hat{b}$ is the estimate of $b_{n-1}$. Because of its explicit physical meaning, modified Gipps’ model was used in several traffic simulation software[22, 23]. Krauss’s model [24] and Gipps’ model are categorized within the same ”family” of CF models, because both models fundamentally consider braking distances to derive interactions between vehicles. This shared focus on braking distances and the resultant dynamics establishes a conceptual link between the two models. Rakha and Wang [25] developed an analytic calibration steps for Gipps car-following for steady state behavior. Some studies [26] showed that the Gipps’ model also allows reproducing the Phantom Traffic Jam, although the model is formulated as a first-order category. A modified Gipps model [27] addressed the potential collision when the most severe braking the following vehicle wishes to undertake is harder than the perceived deceleration rate of lead vehicle.
Newell’s Model In Newell’s Model [28], the time-space trajectory of the $n^{th}$ vehicle is essentially the same as the $(n-1)^{th}$ vehicle, except for a translation in space and in time. This was a very simple rule for car following behavior compared with other models; however, it is often much more accurate than many. Newell’s Model is defined as:
@@
Gipps Model
Gipps (1981) [21]
@@
Newell’s Model
Newell (2002) [28];
@@
1. Limited Representations: Newell’s model simplifies the driving process. Its assumptions about uniform vehicle behavior and simple car-following dynamics may not adequately represent the interactions in complex environments.
Cellular Automata
Wolfram (1992) [37]
@@
GM Models (Gazis-Herman-Rothery)
Chandler et al. (1958) [56]
@@
1. Stimulus-response foundation: The GM model provides generalizable analytic framework from the stimulus-response perspective. $Response=f(sensitivity,stimulus)$. It also allows customization based on different driver characteristics.
1. Deterministic Stimulus: GM model assumes precise perception of stimulus, ignoring the inherent vagueness of CF process. The driver’s reaction is solely based on relative speed, ceasing acceleration or deceleration once relative speed reaches zero, regardless of the distance headway.
@@
Optimal Velocity Model (OVM)
Bando et al. (1995) [65]
@@
Intelligent Driver Model (IDM)
Treiber et al. (2000) [79]
@@ II-A2 Second-Order Kinematic Models
By introducing an internal dynamical variable to represent the subjective level of service, Treiber and Helbing[80] build the IDMM (intelligent driver model with memory) to incorporate memory effects in microscopic traffic models. The Multi-anticipative IDM models [81] consider essential finite reaction times, estimation errors, and spatial and temporal anticipation. Kesting et al. [82] enhanced the IDM model to assess the impact of adaptive cruise control vehicles on traffic capacity. Liebner et al. [83] applied IDM for ADAS to incorporate the spatially varying velocity profile to represent both car-following and turning behavior. Derbel et al. [84] modified the IDM to improve driver safety and respect vehicle capability in emergent situations by adding a new term $cv^{2}/b$ to the expected safe time distance formula. Li et al. [85] modified IDM with power cooperation to strengthen the power of each vehicle in proportion to the immediately preceding vehicle. To build a realistic agent model, Eggert et al. [86] proposed the Foresighted Driver Model (FDM), which assumes that a driver acts in a way that balances predictive risk with utility. Milanés [87] calibrated CACC model and tested IDM on actual vehicles as reference for modeling ACC car-following behavior in traffic flow simulations. Hoermann et al. [88] extended the IDM for autonomous vehicles to describe a probabilistic motion prediction applicable to long-term trajectory planning. Treiber and Kesting [89] added external noise and action points to the IDM to study mechanisms behind traffic flow instabilities and indifferent regions of finite human perception thresholds. Hubmann et al. [90] extend the IDM model for complex urban scenarios by incorporating a behavior prediction framework with context-dependent upper and lower bounds on acceleration. A recent study [91] analyzed the limitations of the IDM model from the numerical analysis: the velocities of specific vehicles might become negative at specific times and might diverge to $\infty$- ∞ in a finite time. The Self-Adaptive IDM (SA-IDM) [92] dynamically adjusts acceleration in response to the leading vehicle’s speed and time headway, ensuring stable and overdamped platoon behavior under all conditions.

### v5

@@ Abstract
Car-following (CF) algorithms are crucial components of traffic simulations and have been integrated into many production vehicles equipped with Advanced Driving Assistance Systems (ADAS). Insights from the model of car-following behavior help researchers to understand the causes of various macro phenomena that arise from interactions between pairs of vehicles. Car-following Models encompass multiple disciplines, including traffic engineering, physics, dynamic system control, cognitive science, machine learning, deep learning, and reinforcement learning. This paper presents an extensive survey highlighting the differences, complementarities, and overlaps among microscopic traffic flow and control models based on their underlying principles and design logic. It reviews a range of representative algorithms, from theory-based Kinematic Models, Psycho-Physical Models, and Adaptive Cruise Control Models to learning-based algorithms like Reinforcement Learning (RL) and Imitation Learning (IL). Additionally, it considers the impact of large Generative AI (GenAI) models, categorized as Knowledge-Driven models. The survey discusses the strengths and limitations of these models, explores their applications in various contexts, and summarizes available datasets across different domains to address knowledge gaps. By synthesizing historical developments and current trends, this survey provides a comprehensive overview of the evolution and future directions of car-following models. It highlights the importance of interdisciplinary approaches and the need for continued innovation to address emerging challenges.
@@
Car-Following Behavior, Adaptive Cruise Control, Theory-based, Learning-based, Knowledge-driven.
@@ I Introduction
To furnish a complete picture of traffic flow, vehicles on the road are studied as a whole - not merely as isolated particles. Car-following (CF) is the fundamental driving behavior that forcefully constrains the ego-vehicle from adopting a specific pattern in the presence of other vehicles (Figure 1). Since the 1950s or even earlier, many researchers have focused on driver behavior modeling and applied such models to the simulation of Human Driver Vehicles (HDVs) to investigate micro- and macro-traffic phenomena [1, 2]. The expansion of autonomy and connectivity in the transportation system has escalated the hype for integration of car following models with Connected and Automated Vehicle (CAV) technologies, such as collaborative driving, and Vehicle-to-Everything (V2X) communication. Autonomous Vehicles (AVs) with different parameter settings or objective functions become a new source of variety and add more complexities to the existing roadway systems. The heterogeneity of vehicle types and driver behaviors gives rise to various challenges for traffic operation, which must be comprehensively analyzed before upgrading existing infrastructure. Car-following models become a powerful tool to address challenges mentioned above and answer those critical questions in the upcoming era of the connected and automated transportation systems. By analyzing the interaction between vehicles, researchers can identify efficient driving strategies that includes optimizing speeds, minimizing unnecessary acceleration and deceleration events, and maintaining optimal following distances to ensure smooth traffic flow. Modern adaptive cruise control (ACC) systems already use basic car-following principles to maintain safe distances between vehicles.
@@
Existing survey papers[3, 4, 5, 6, 7, 8, 9, 12, 10, 11, 13, 14, 15, 16, 17, 18, 19] lack of structured framework and often have limitations in terms of scope and perspective (see Table I). For example, a frequently referenced paper [4] mainly surveyed from traffic psychologic perspective and did not consider Intelligent Driver Model (IDM) and Optimal Velocity Model (OVM). Some surveys [11, 9] reviewed multiscale traffic flow models from the engineering and physics perspective, while many control theory models were omitted. Other literature reviews [8, 10] are limited to theory-based models and inadequate to reflect AI/ML models. By taking into account various domains and angles, this review paper aims to provide a more structured categorization of the car following models (see Figure 2), addressing the potential gaps left by previous surveys with a narrower focus. A multidisciplinary review can be particularly beneficial in fields like intelligent vehicle technology, where insights from different disciplines – such as electrical & computer engineering, mechanic engineering, traffic flow theory, physics, and optimal control – are crucial for developing a complete picture of how these agents operate and interact with their environment. This review paper not only integrates diverse knowledge but also promotes the discovery of innovative solutions that might be overlooked when considering a single perspective. Our contributions can be itemized as follows:
  • Multidisciplinary Perspective: Unlike previous surveys that focus on a single aspect of car-following models (e.g., kinematic characteristics, specific applications, or human factors), our paper provides a comprehensive review from multiple disciplinary perspectives. This broader approach allows for a more thorough evaluation of the state-of-the-art, challenges, and future directions of car-following models, making a unique contribution to the field.
  • Novel Taxonomy for Car-Following Models: We introduce an effective taxonomy to categorize the broad spectrum of car-following model research. Our taxonomy is not just a listing; it delves into the underlying design principles, offering a structured framework that surpasses the categorizations found in previous surveys. This novel approach allows for a more intuitive understanding of the field’s evolution and current state.
  • Extensive Coverage of Application Cases: Our survey extends the discussion to various application cases built upon car-following models, highlighting their significant impact on the environment, safety, mobility, and operation efficiency. This comprehensive review of applications is unprecedented in existing literature, providing valuable insights into the practical implications of car-following models.
  • Future Directions with Human-like Reasoning: We underscore the future direction of integrating AI-driven methods with theory-based approaches to embed human-like reasoning processes into car-following models. This perspective is critical for developing more advanced, realistic models that can better mimic human driving behaviors, addressing a gap in current research trends.
By providing a thorough evaluation of current challenges and paving the way for future research directions in car-following models, our review marks a distinctive and valuable contribution to the literature.
@@
Car following models [3] 1992 Summarized car-following models using mathematical model descriptions, stability analysis, calibration from empirical testing and evaluation and finally automatic vehicular control systems. It only covers stimulus-response models from psychological and behavior perspective.
@@
Car-following: a historical review [4] 1999 Summarized theoretical early-stage car-following models from psychological and behavior perspective, such as safety distance, desired gaps, stimulus-response process. It doesn’t include some important models such as Intelligent Driver Model (IDM) and Optical Velocity Model (OVM).
@@
Comparison of Car-following models [5] 2004 Compared theoretical car-following models used in well-known traffic micro-simulation software, including AIMSUN, MITSIM, Paramics and VISSIM. From a software perspective to evaluate how different programs could replicate real world scenarios.
@@
Driving Behavior: models and challenges [6] 2007 Reviewed and expanded review lists in [4] with IDM, OVM and Cellular Automata. Discussed the challenges associated with the myopic driver behavior models. Only listed theory-based car-following models from human behavior perspective.
Car-following models fifty years of linear stability analysis a mathematical perspective  [7] 2011 A comprehensive survey of linear stability types, illustrated with straightforward examples and rigorous mathematical criteria. From a control theory perspective, it introduced a general framework that simplifies the analysis of linear stability by requiring only three partial derivatives.
@@
Microscopic car-following model for the traffic flow: the state of the art [8] 2012 Summarized and reviewed theory-based car-following models, also provided the stability analysis from control theory point of view. Only reviewed theory-based car-following models from human behavior perspective.
@@
Traffic flow dynamics[9] 2013 Categorized existing models as elementary models and strategic models by design principles in microscopic traffic flow modeling. Only covers kinematic models from the physics perspectives.
@@
Incorporating human-factors in car-following models: a review of recent developments and research needs[10] 2014 Introduced various human factors that could potentially impact driving styles. Only summarized theory-based car-following models from engineering and human behavior perspective.
@@
Genealogy of traffic flow models [11] 2015 Summarized macro-, meso- and micro- traffic flow models, offered an historical model tree by tracing genealogy of the models. Reviewed and surveyed car following models from the traffic engineering and physical perspectives
@@
A state-of-the-art review of car-following models with particular considerations of heavy vehicles [12] 2015 Categorized CF models into classical models and AI models. Emphasized the impact of heavy trucks in car following behavior. Under AI category, only fuzzy logic models and Neural Network models were considered. Under the classical model category, it doesn’t include most recent ACC models
@@
Simulation Strategies for Mixed Traffic Conditions: A Review of Car-Following Models and Simulation Frameworks [13] 2020 This review paper focus on Heterogeneous vehicle types and mixed traffic flow modeling Inherit the same categorization from previous review paper [4] and mainly from simulation software perspective
@@
A review of car-following models and modeling tools for human and autonomous-ready driving behaviors in micro-simulation [14] 2021 Emphasized the differences between human-driven vehicles and autonomous vehicles in car-following models. Inherit the same categorization from previous review paper [4] and mainly from simulation software perspective.
@@
The Car-Following Model and Its Applications in the V2X Environment: A Historical Review [15] 2021 Emphasize applications of car-following models in the V2X environment Use the similar categorization framework from previous review paper [4, 6] to categorize car-following models
@@
Modeling the Car-Following Behavior with Consideration of Driver, Vehicle, and Environment Factors: A Historical Review [16] 2022 This review paper surveyed car models that consider diverse driver styles, physical and dynamic characteristics of vehicles, as well as environmental factors. Inherit the similar categorization from previous review paper [4, 6, 15] and is developed from traffic engineering perspective.
Modelling and simulation of (connected) autonomous vehicles longitudinal driving behavior: A state-of-the-art [17] 2023 This paper reviewed parameters for (C)AVs modelling. It also summarizes the lately published data-driven models for (C)AVs CF behavior, and identification of KPIs used for impact assessments. Mainly focus on CF models’ parameters that are more crucial and sensitive in differentiating (C)AVs from human driven vehicles.
Microscopic Modelling of Car-Following Behavior: Developments and Future Directions [18] 2023 This paper analyzes advantages and limitations of four main types of CF models: Kinematics-based, dynamics-based, psychological-based, and learning-based. Didn’t provide subcategories and missed Cellular Automata, and ACC models.
Car-following models for human-driven vehicles and autonomous vehicles: A systematic review [19] 2023 Reviewed CF models in terms of accuracy, calibrations, stability analysis and specific model parameters for HDVs and AVs simulation Limited to Traffic Engineering perspective for categorization and applicable background
@@
Figure 2: Proposed Taxonomy of Car Following Models (CFMs) and Representative Algorithms. Kinematic Models are divided into first-order and second-order subcategories. Knowledge-driven Large GenAI Models are a revolutionary category, providing new insights and solutions with commonsense.
@@ II-A1 First-Order Kinematic Models
In the formulation, $a_{n}$ is the maximum acceleration; $b_{n}$ is the most severe braking that the driver of vehicle $n$ wishes to undertake; $v_{n}$ is the speed at which the driver of vehicle $n$ wishes to travel; $V_{n}$ is the desired speed of vehicle $n$; $\tau$ is the reaction time; $\hat{b}$ is the estimate of $b_{n-1}$. Because of its explicit physical meaning, modified Gipps’ model was used in several traffic simulation software[26, 27]. Krauss’s model [28] and Gipps’ model are categorized within the same ”family” of CF models, because both models fundamentally consider braking distances to derive interactions between vehicles. This shared focus on braking distances and the resultant dynamics establishes a conceptual link between the two models. Rakha and Wang [29] developed an analytic calibration steps of Gipps model for steady state behavior. Some studies [30] showed that the Gipps’ model also allows reproducing the Phantom Traffic Jam, although the model is formulated as a first-order category. A modified Gipps model [31] addressed the potential collision when the most severe braking the following vehicle wishes to undertake is harder than the perceived deceleration rate of lead vehicle.
Newell’s Model In Newell’s Model [32], the time-space trajectory of the $n^{th}$ vehicle is essentially the same as the $(n-1)^{th}$ vehicle, except for a translation in space and in time. This is a very simple rule for car following behavior compared with other models; however, it is often much more accurate than many. Newell’s Model is defined as:
@@
Gipps Model Gipps (1981) [25]
@@
Newell’s Model Newell (2002) [32];
@@
1. Simplicity: Newell’s Model is easy to implement and analyze.
2. Traffic wave simulation: Newell’s Model is often used to simulate traffic waves caused by traffic signals or other disturbances in the interrupted traffic flow.
1. Limited Representations: Newell’s Model simplifies the driving process without any driver-related parameters, which may lead to less realistic representations.
Cellular Automata Wolfram (1992) [41]
@@
GM Models (Gazis-Herman-Rothery) Chandler et al. (1958) [60]
@@
1. Stimulus-response foundation: The GM model provides generalizable analytic framework from the stimulus-response perspective. $Response=f(sensitivity,stimulus)$. It also allows customization based on different driver characteristics. 1. Deterministic Stimulus: GM model assumes precise perception of stimulus, ignoring the inherent vagueness of CF process.
@@
Optimal Velocity Model (OVM) Bando et al. (1995) [69]
@@
Intelligent Driver Model (IDM) Treiber et al. (2000) [83]
@@ II-A2 Second-Order Kinematic Models
By introducing an internal dynamical variable to represent the subjective level of service, Treiber and Helbing[84] build the IDMM (intelligent driver model with memory) to incorporate memory effects in microscopic traffic models. The Multi-anticipative IDM models [85] consider essential finite reaction times, estimation errors, and spatial and temporal anticipation. Kesting et al. [86] enhanced the IDM model to assess the impact of adaptive cruise control vehicles on traffic capacity. Liebner et al. [87] applied IDM for ADAS to incorporate the spatially varying velocity profile to represent both car-following and turning behavior. Derbel et al. [88] modified the IDM to improve driver safety and respect vehicle capability in emergent situations by adding a new term $cv^{2}/b$ to the expected safe time distance formula. Li et al. [89] modified IDM with power cooperation to strengthen the power of each vehicle in proportion to the immediately preceding vehicle. To build a realistic agent model, Eggert et al. [90] proposed the Foresighted Driver Model (FDM), which assumes that a driver acts in a way that balances predictive risk with utility. Milanés [91] calibrated CACC model and tested IDM on actual vehicles as reference for modeling ACC car-following behavior in traffic flow simulations. Hoermann et al. [92] extended the IDM for autonomous vehicles to describe a probabilistic motion prediction applicable to long-term trajectory planning. Treiber and Kesting [93] added external noise and action points to the IDM to study mechanisms behind traffic flow instabilities and indifferent regions of finite human perception thresholds. Zhou, Qu and Jin (2017) [94] added multiplication factors to build Cooperative Intelligent Driver Model(CIDM) to simulate AVs behavior to show that AVs could relieve traffic oscillation by give larger gaps for on-ramp vehicles. Hubmann et al. [95] extend the IDM model for complex urban scenarios by incorporating a behavior prediction framework with context-dependent upper and lower bounds on acceleration. A recent study [96] analyzed the limitations of the IDM model from the numerical analysis: the velocities of specific vehicles might become negative at specific times and might diverge to $\infty$- ∞ in a finite time. To capture the stochastic nature of human driver behavior, IDM with time-varying parameters is proposed, with each parameter as a stochastic process and calibrating them in a time-varying manner [97]. The Self-Adaptive IDM (SA-IDM) [98] dynamically adjusts acceleration in response to the leading vehicle’s speed and time headway, ensuring stable and overdamped platoon behavior under all conditions.